Black Body Radiation:

A 100W incandescent bulb has a tungsten filament at 2000C. If you model the bulb as a blackbody where the average photon energy is given by the peak wavelength of the blackbody spectrum, how many photons N is the bulb radiating per second?
Answer #1

Here ,

temperature of bulb , T = 2000 degree C

T = 2273 K

now , wavelength is

wavelength = 2.898 *10^-3/T

wavelength = 2.898 *10^-3/2273

wavelength = 1.275 *10^-6 m

energy of 1 photon , E = h*c/wavelength

E = 6.626 *10^-34 * 3 *10^8/( 1.275 *10^-6)

E = 1.559 *10^-19 J

total number of photon = 100/( 1.559 *10^-19)

total number of photon = 6.414 *10^20

the total number of photon is 6.414 *10^20